There is a necessary gap of $500 million in real output. The marginal propensity to save is .125. If a change in tax revenue is prescribed by the government, how much should the government decrease tax revenue?
Answer #1

Tax multiplier = - (1 - MPS) / MPS = - (1 - 0.125) / 0.125 = - 0.875 / 0.125 = - 7

Required decrease in tax = 500 m / 7 = 71.4285 m

= 71.43 m